### v1

# A Quantile Nelson-Siegel model

Abstract
A widespread approach to modelling the interaction between macroeconomic variables and the yield curve relies on three latent factors usually interpreted as the level, slope, and curvature (Diebold et al., 2006). This approach is inherently focused on the conditional mean of the yields and postulates a dynamic linear model where the latent factors smoothly change over time.
However, periods of deep crisis, such as the Great Recession and the recent pandemic, have highlighted the importance of statistical models that account for asymmetric shocks and are able to forecast the tails of a variable’s distribution. A new version of the dynamic three-factor model is proposed to address this issue based on quantile regressions. The novel approach leverages the potential of quantile regression to model the entire (conditional) distribution of the yields instead of restricting to its mean.
An application to US data from the 1970s shows the significant heterogeneity of the interactions between financial and macroeconomic variables across different quantiles. Moreover, an out-of-sample forecasting exercise showcases the proposed method’s advantages in predicting the yield distribution tails compared to the standard conditional mean model. Finally, by inspecting the posterior distribution of the three factors during the recent major crises, new evidence is found that supports the greater and longer-lasting negative impact of the great recession on the yields compared to the COVID-19 pandemic.
Keywords: Bayesian inference; Nelson-Siegel model; Quantile regression; Time-varying parameters; Yield curve.
1 Introduction
The ongoing global challenges stemming from the COVID-19 pandemic and recent geopolitical conflicts have prompted a renewed focus on inflation and monetary policy. Of particular significance is the information embedded in the term structure of interest rates, offering valuable insights into the evolution of these economic factors. The slope of the yield curve has gained considerable recognition as a key predictor of economic activity, exerting a substantial influence on investors’ decision-making processes.
This significance is underscored by research from Benzoni et al. (2018) and Haubrich (2021), which highlights a strong correlation between an inversion of the yield curve and the onset of recessions. For instance, in May 2019, the yield curve inverted almost a year before the commencement of the most recent recession in March 2020. More recently, Fonseca et al. (2023) emphasised the unprecedented inversion in the risk-free yield curves in the United States due to the rapid rise in short-term interest rates since 2022. They pointed out that “recession probability models based solely on the yield curve slope currently point to elevated odds of recession in the euro area and the United States in one year’s time” thereby stressing the critical importance of dealing with and studying the yield curve in contemporary macroeconomic analysis.
Transitioning to a macroeconomic perspective, significant time variations have been shown to influence the dynamics of US inflation and real activity (Mumtaz and Surico, 2009). Mönch (2012) studied the evolution of macroeconomic variables by using level, slope, and curvature factors, finding evidence of the informativeness of the curvature factor about the future evolution of the yield curve. Moving to factor analysis, Coroneo et al. (2016) provided evidence that some common factors drive the dynamics of macroeconomic variables and government bonds. More recently, Fernandes and Vieira (2019) predicted the yield curve using different forward-looking macroeconomic variables and highlighted the importance of these variables in forecasting the yield curve.
In the context of modelling the term structure of interest rates, the Nelson-Siegel (NS) model, introduced by Nelson and Siegel (1987), is the most popular empirical framework employed within the literature. Diebold and Li (2006) generalised this model by introducing smooth dynamics for the latent factors, resulting in the dynamic NS (DNS) model, which outperforms the standard NS approach in an out-of-sample forecasting context. However, these approaches do not account for the potential impact of macroeconomic variables on the yield curve. To address this limitation, Diebold et al. (2006) and Coroneo et al. (2016) analysed the dynamic linkages between yield curve factors and macroeconomic factors in the US. Koopman and van der Wel (2013) found evidence of interdependence between macroeconomic and latent factors in the US term structure, thus supporting the use of macroeconomic variables in combination with financial yield curve factors. Continuing along this line of research, Bianchi et al. (2009) applied a macro-factors augmented DNS model with time-varying parameters to UK data. Recently, Han et al. (2021) introduced a time-varying decay parameter for the arbitrage-free NS model and demonstrated that the relative factor loading reaches its peak just before starting the decline right before a recession.
It is worth emphasising that the literature on the term structure and the analysis of the relationship between yield curve factors and macroeconomic variables so far has been concerned with conditional mean models. However, the in-sample and forecasting performances of this approach are only sufficient in relation to the mean of the yields, as the method is not suited to investigate the complete distribution of the yields. To address the limitations of standard conditional mean Nelson-Siegel models and capture the complete picture of the conditional distribution of the yields, we propose a novel quantile regression Nelson-Siegel model with time-varying parameters (TVP-QR-NS) model.
We contribute to this literature by proposing a quantile version of the Nelson-Siegel model for the yield curve. Specifically, we extend the Diebold et al. (2006) ’Yields-Macro’ model to a quantile framework. Our approach assumes that the conditional $\tau$-quantile of the yields is given by the Nelson-Siegel three-factor specification, where the factors are (smoothly) time-varying. Quantile regression (QR, see Koenker and Bassett, 1978) does not assume a parametric likelihood for the conditional distribution of the response variable. Unlike standard models based on the conditional mean and relying on symmetric second-moment dynamics, our novel framework offers robust modelling of the conditional quantiles, thus enabling a comprehensive investigation of the entire conditional distribution. Moreover, by considering multiple quantile levels, we can follow the method proposed by Mitchell et al. (2022) and combine the quantiles to obtain the distributional yield curve at each point in time, as shown in Figure 1. This permits the investigation of dynamic asymmetry within the distribution of the term structure of interest rates. This feature is particularly crucial when undertaking the modelling and forecasting of interest rate term structures, as emphasised in our empirical application.
Figure 1: Distributional term structure of the interest rate for June and December 2022.
Inference is performed within a Bayesian framework by leveraging the connection between a likelihood formed by the asymmetric Laplace distribution (Kotz et al., 2001) and quantile regression provided by Yu and Moyeed (2001), and Petrella and Raponi (2019).
The main feature of the proposed TVP-QR-NS model is the coupling of quantile regression with time-varying three-factor Nelson-Siegel specification. In addition, to cope with the potentially high dimensionality of the latent dynamic factors, we use a precision sampler (Chan and Jeliazkov, 2009; Chan et al., 2020) that allows us to sample the entire path of the factors jointly without any loop. Overall, this combination provides a simple framework and computationally efficient method to investigate and forecast the term structure at every quantile level of interest.
The proposed method is then applied to investigate and forecast the US term structure of interest rates. Specifically, an in-sample analysis highlights the advantages of the quantile approach compared to the standard conditional mean Nelson-Siegel model. In particular, we are able to obtain the entire distribution of the three factors at different points in time, thus enabling a more detailed analysis of the effects of the recent crises (the Great Recession and the COVID-19 pandemic) on the term structure. Interestingly, these results, which could not be obtained using a conditional mean approach, reveal substantial differences between the impacts of the financial crisis and COVID-19. The former shifted leftwards the distribution of the short-term interest rate without meaningful impact on the long-term. Conversely, the pandemic had an initial significant adverse effect on both the short- and long-term rates, but the distribution in 2022 shifted rightwards compared to the pre-COVID one.
The advantages of the proposed TVP-QR-NS model are better highlighted in terms of tail risk yield forecasting. Compared to the conditional mean model, our approach provides better out-of-sample forecasts in both tails of the yield distribution, as measured by the quantile score. The difference is particularly striking on the left tail and consistent over time.
The remainder of this article is organised as follows. Section 2 introduces and describes the novel framework, then Section 3 presents the Bayesian approach to inference. Section 4 details the empirical application. Finally, Section 5 concludes.
2 Methodology
2.1 Nelson-Siegel Yield Curve model
The standard Nelson-Siegel (NS) model of the yield curve assumes that
$y_{t}(m)=\beta_{1}+\beta_{2}\left(\frac{1-\exp^{-\lambda m}}{\lambda m}\right)+\beta_{3}\left(\frac{1-\exp^{-\lambda m}}{\lambda m}-\exp^{-\lambda m}\right),$ (1)
where $m$ denotes the time to maturity; that is, given a set of parameters, the yield curve is a deterministic function of maturity. Diebold and Li (2006) pioneered the stochastic modelling of the yield curve and suggested the linear model
$y_{t}(m)=\beta_{1t}+\beta_{2t}\left(\frac{1-\exp^{-\lambda m}}{\lambda m}\right)+\beta_{3t}\left(\frac{1-\exp^{-\lambda m}}{\lambda m}-\exp^{-\lambda m}\right)+\epsilon_{t}(m)$ (2)
where $\epsilon_{t}(m)\sim\mathcal{N}(0,\sigma^{2}(m))$ is an independent normal innovation. Therefore, given a set of time-varying parameters, the expected value of the curve is a deterministic function of maturity.
2.2 A Time-Varying Parameters Quantile NS Model
Assume that $y_{t}(m)\in\mathcal{Y}\subset\mathds{R}$ for all maturities $m\in\mathds{R}_{+}$, and let $F_{t}(\cdot;m)$ denote the cumulative distribution function of $y_{t}(m)$. We propose a new approach that alleviates the parametric assumptions of $y_{t}(m)$’s by modelling the quantile function of $F$,
$Q_{t}^{\tau}(m)=\inf\left\{x\in\mathcal{Y}:F_{t}(x;m)\geq\tau\right\}$
for each quantile level $\tau\in(0,1)$ as follows
$Q_{t}^{\tau}(m)=\beta_{1t}(\tau)+\beta_{2t}(\tau)\Big{(}\frac{1-e^{-\lambda^{\tau}m}}{\lambda^{\tau}m}\Big{)}+\beta_{3t}(\tau)\Big{(}\frac{1-e^{-\lambda^{\tau}m}}{\lambda^{\tau}m}-e^{-\lambda^{\tau}m}\Big{)},$ (3)
where $m$ is the maturity, $\beta_{t1}^{\tau}$, $\beta_{t2}^{\tau}$, and $\beta_{t3}^{\tau}$ denote the (unobserved) level, slope and curvature factors, and $\lambda^{\tau}\in\mathds{R}_{+}$ is a decay parameter. In the following, we shall assume $\lambda^{\tau}=\lambda$. Thus, by directly modelling the quantile function, we allow the yield curve factors to behave differently across the quantile levels.
Consider a series of observations at specific maturities $m\in\{1,\ldots,M\}$ and denote with $\mathbf{y}_{t}=(y_{t}(m_{1}),\ldots,y_{t}(m_{M}))^{\prime}$ the vector of yields at different maturity. Let $\mathbf{x}(m;\lambda)=\big{(}1,\frac{1-e^{-\lambda m}}{\lambda m},\frac{1-e^{-\lambda m}}{\lambda m}-e^{-\lambda m}\big{)}^{\prime}$, and $\mathbf{X}(\lambda)=\left[\mathbf{x}(m_{1};\lambda),\mathbf{x}(m_{2};\lambda),\ldots,\mathbf{x}(m_{M};\lambda)\right]^{\prime}\in\mathds{R}^{M\times 3}$. Then, the conditional quantile model can be written as
$Q^{\tau}_{t}=\mathbf{X}(\lambda)\bm{\beta}_{t}(\tau)$ (4)
such that $Q^{\tau}_{t}=(Q^{\tau}_{t}(m_{1}),\ldots,Q^{\tau}_{t}(m_{M}))^{\prime}$ and $\bm{\beta}_{t}(\tau)=(\beta_{1t}(\tau),\beta_{2t}(\tau),\beta_{3t}(\tau))^{\prime}$. Moreover, at each quantile level $\tau$, the unobserved factors are modelled jointly with a vector of macroeconomic (observed) variables $\mathbf{z}_{t}\in\mathds{R}^{r}$. Specifically, the vector of time-varying (unobserved) yield curve and (observed) macroeconomic factors $\bm{\phi}_{t}=(\bm{\beta}_{t}^{\prime}(\tau),\mathbf{z}_{t}^{\prime})^{\prime}$ is assumed to follow a stationary VAR(1) process
$\bm{\phi}_{t}=\bm{\gamma}_{0}+\Gamma\bm{\phi}_{t-1}+\bm{\eta}_{t},\qquad\bm{\eta}_{t}\sim\mathcal{N}(\mathbf{0},\Omega).$ (5)
Differently from Coroneo et al. (2016) that proposed an approach related to ours in the context of a conditional mean model for the yields, we are modelling the conditional quantile of $y_{t}(m)$ and rely on an alternative specification of the interactions between the yield and the macroeconomic factors.
To design a regression model for the yield such that the conditional $\tau$-quantile corresponds to eq. (3), we adopt a likelihood-based approach and rely on the properties of the asymmetric Laplace distribution (Kotz et al., 2001). In particular, for a given quantile $\tau$, we consider a regression model where the innovation follows an asymmetric Laplace distribution $\epsilon_{t}(m)\sim\mathcal{AL}(\mu,\sigma,\tau)$, with location $\mu\in\mathds{R}$, scale $\sigma>0$, skewness parameter $\tau\in(0,1)$, with density
$p(x|\mu,\sigma,\tau)=\frac{\tau(1-\tau)}{\sigma}\exp\Big{\{}-\rho_{\tau}\Big{(}\frac{x-\mu}{\sigma}\Big{)}\Big{\}},$
where $\rho_{\tau}(x)=x(\tau-\mathbb{I}(x\leq\tau))$ is the check-loss function. Therefore, combining the conditional quantile in eq. (3) with the asymmetric Laplace innovation results in
$y_{t}(m)=\mathbf{x}(m;\lambda)^{\prime}\bm{\beta}_{t}(\tau)+\epsilon_{t}(m),\qquad\epsilon_{t}(m)\sim\mathcal{AL}(0,\sigma,\tau).$ (6)
It is important to highlight that the asymmetric Laplace distribution admits a stochastic representation as a location-scale mixture of Gaussian (Kotz et al., 2001), thus allowing to rewrite eq. (6) as
$y_{t}(m)\stackrel{{\scriptstyle d}}{{=}}\mathbf{x}(m;\lambda)^{\prime}\bm{\beta}(\tau)+w_{m,t}\sigma\theta_{\tau,1}+\sqrt{w_{m,t}}\sigma\theta_{\tau,2}u_{t}(m),\qquad u_{t}(m)\sim\mathcal{N}(0,1),$ (7)
where $w_{m,t}\sim\mathcal{E}xp(1)$ is exponential with rate one. To establish an equivalence between the mode of the asymmetric Laplace distribution and the optimisation of the standard quantile regression problem in Koenker and Bassett (1978), the parameters $\theta_{\tau,1},\theta_{\tau,2}$ are constrained to the following values (Kotz et al., 2001):
$\theta_{\tau,1}=\frac{1-2\tau}{\tau(1-\tau)},\qquad\theta_{\tau,2}=\sqrt{\frac{2}{\tau(1-\tau)}}.$ (8)
Throughout the article, we follow the existing literature and fix the decay parameter at $\lambda=0.0609$ (e.g., see Diebold and Li, 2006). Moreover, the scale of the innovation is set to $\sigma=1.0$, as in the original quantile regression model (Yu and Moyeed, 2001).
The combination of eq. (5) and (6) provides the state-space form of the proposed time-varying parameters quantile Nelson-Siegel model (TVP-QR-NS) as
$\begin{split}y_{t}(m)&=\mathbf{x}(m;\lambda)^{\prime}\bm{\beta}_{t}(\tau)+\epsilon_{t}(m),\qquad\epsilon_{t}(m)\sim\mathcal{AL}(0,\sigma,\tau)\\ \bm{\phi}_{t}&=\bm{\gamma}_{0}+\Gamma\bm{\phi}_{t-1}+\bm{\eta}_{t},\qquad\bm{\eta}_{t}\sim\mathcal{N}(\mathbf{0},\Omega),\end{split}$ (9)
where we recall the fixed values $\lambda=0.0609$ and $\sigma=1.0$. Denoting with $\mathbf{Y}=(\mathbf{y}_{1},\ldots,\mathbf{y}_{T})^{\prime}\in\mathds{R}^{T\times M}$ the matrix of observations, the joint density of the observables and auxiliary variables (conditional on the time-varying factors) is
$p(\mathbf{Y},\mathbf{w}|\bm{\beta})$ $=\prod_{m=1}^{M}\mathcal{E}xp(w_{m,t}|1)\times\mathcal{N}\big{(}y_{t}(m)|\mathbf{x}(m;\lambda)^{\prime}\bm{\beta}_{t}(\tau)+w_{m,t}\sigma\theta_{\tau,1},\>w_{m,t}\sigma^{2}\theta_{\tau,2}^{2}\big{)}.$ (10)
3 Estimation
This section is devoted to the prior description and the derivation of the precision sampler algorithm.
3.1 Prior Specification
We consider a multivariate Gaussian prior for the initial value of the time-varying factors:
$\bm{\beta}_{0}(\tau)\sim\mathcal{N}(\underline{\bm{\mu}}_{\beta},\underline{\Sigma}_{\beta}).$ (11)
By leveraging the mixture representation of the asymmetric Laplace distribution, we obtain a conditionally linear Gaussian state space for the latent dynamic vector $\bm{\beta}_{t}$. Thus, we sample from the posterior full conditional distribution of $\bm{\beta}_{t}$ directly in a single loop, using a precision sampler (e.g., see Chan and Jeliazkov, 2009; Chan et al., 2020) as described in the following subsection.
Let us define $e_{m,t}=y_{t}(m)-\mathbf{x}(m;\lambda)^{\prime}\bm{\beta}_{t}$. Then, the posterior full conditional distribution of the auxiliary variable $w_{t}$ is obtained as
$w_{m,t}|y_{t}(m),\bm{\beta}_{t}$ $\propto w_{m,t}^{-\frac{1}{2}}\exp\Big{\{}-\frac{1}{2}\Big{[}\frac{\theta_{\tau,1}^{2}}{\theta_{\tau,2}^{2}}w_{m,t}+2w_{m,t}+w_{m,t}^{-1}\frac{e_{m,t}^{2}}{\theta_{\tau,2}^{2}\sigma^{2}}\Big{]}\Big{\}},$ (12)
which is the kernel of the generalized inverse Gaussian distribution $\text{GiG}(\overline{p}_{w},\overline{a}_{w},\overline{b}_{w,m,t})$ with
$\overline{p}_{w}=1-\frac{1}{2},\qquad\overline{a}_{w}$ $=2+\frac{\theta_{\tau,1}^{2}}{\theta_{\tau,2}^{2}},\qquad\overline{b}_{w,m,t}=\frac{e_{m,t}^{2}}{\theta_{\tau,2}^{2}\sigma^{2}}.$
Finally, we specify a Gaussian prior for the vectorised coefficient matrix of the state equation, $\bm{\gamma}=(\bm{\gamma}_{0}^{\prime},\operatorname{vec}(\Gamma)^{\prime})^{\prime}$, and an inverse Wishart prior for the innovation covariance, that is
$\bm{\gamma}\sim\mathcal{N}(\underline{\bm{\mu}}_{\Gamma},\underline{\Sigma}_{\Gamma}),\qquad\Omega\sim\mathcal{IW}(\underline{v},\underline{V}).$ (13)
As the prior distributions are conjugate, the full conditional posteriors are straightforward to obtain as
$\bm{\gamma}|\bm{\phi},\Omega\sim\mathcal{N}(\overline{\bm{\mu}}_{\Gamma},\overline{\Sigma}_{\Gamma}),\qquad\Omega|\bm{\phi},\gamma_{0},\Gamma\sim\mathcal{IW}(\overline{v},\overline{V}).$ (14)
The derivation of these full conditional posteriors is standard (e.g, see Chan, 2020).
3.2 The precision sampler
While the conditional distribution of the parameters given the data and the set of unobserved latent factors is relatively standard, the most significant computational costs arise from sampling the large number of latent factors, denoted as $\bm{\beta}_{t}(\tau)$.
Let us start by rewriting the measurement equation in (9) in compact form by stacking all observations together. Specifically, we define the stacked vectors $\mathbf{y}=(\mathbf{y}_{1}^{\prime},\ldots,\mathbf{y}_{T}^{\prime})^{\prime}$, $\bm{\beta}=(\bm{\beta}_{1}(\tau)^{\prime},\ldots,\bm{\beta}_{T}(\tau)^{\prime})^{\prime}$, $\mathbf{w}=(w_{1},\ldots,w_{T})^{\prime}$, and $\mathbf{u}=(u_{1},\ldots,u_{T})^{\prime}$, we have
$\mathbf{y}=\mathbf{I}_{T}\otimes\mathbf{X}(\lambda)\bm{\beta}+\theta_{\tau,1}\sigma\mathbf{w}+\mathbf{u},$ (15)
where $\mathbf{u}\sim\mathcal{N}(\mathbf{0}_{TM},\sigma^{2}\theta_{\tau,2}\Omega_{1})$ with $\Omega_{1}=\operatorname{diag}(\mathbf{w})$. Similarly, we define $\bm{\phi}=(\bm{\phi}_{1}^{\prime},\ldots,\bm{\phi}_{T}^{\prime})^{\prime}$ and rewrite the state equation in a matrix form as
$\mathbf{H}\bm{\phi}=\tilde{\bm{\gamma}}+\bm{\eta},\qquad\bm{\eta}\sim\mathcal{N}\left(\mathbf{0}_{T(r+3)},\Omega\otimes\mathbf{I}_{T}\right),$ (16)
where
$\mathbf{H}=\begin{pmatrix}\mathbf{I}_{r+3}&\mathbf{0}_{r+3}&\mathbf{0}_{r+3}&\mathbf{0}_{r+3}&...&\mathbf{0}_{r+3}&\mathbf{0}_{r+3}\\ \mathbf{0}_{r+3}&-\Gamma&\mathbf{I}_{r+3}&\mathbf{0}_{r+3}&...&\mathbf{0}_{r+3}&\mathbf{0}_{r+3}\\ ...&...&...&...&...&...&...\\ ...&...&...&...&...&\mathbf{I}_{r+3}&\mathbf{0}_{r+3}\\ ...&...&...&...&...&-\Gamma&\mathbf{I}_{r+3}\end{pmatrix},\qquad\tilde{\bm{\gamma}}=\begin{pmatrix}\bm{\gamma}_{0}\\ \bm{\gamma}_{0}\\ ...\\ \bm{\gamma}_{0}\\ \bm{\gamma}_{0}\end{pmatrix}.$
Note that the unobserved factors are placed first in the VAR vector such that
$\bm{\phi}=S_{1}\bm{\beta}+S_{2}\mathbf{u},$
where
$S_{1}=\mathbf{I}_{T}\otimes(\mathbf{I}_{3},\mathbf{0}_{3\times r})^{\prime},\quad S_{2}=\mathbf{I}_{T}\otimes(\mathbf{0}_{3\times r},\mathbf{I}_{r})^{\prime}.$
Thus, we obtain the conditional posterior of the time-varying yield curve factors as $\bm{\beta}|\mathbf{Y},\mathbf{w},\Gamma,\bm{\gamma}_{0},\Omega\propto\mathcal{N}\left(\bm{\mu}_{b},K_{b}^{-1}\right)$, where
$K_{b}$ $=\frac{1}{\sigma^{2}\theta_{\tau,2}}\left(\mathbf{I}_{T}\otimes\mathbf{X}(\lambda)\right)^{\prime}\Omega_{1}^{-1}\left(\mathbf{I}_{T}\otimes\mathbf{X}(\lambda)\right)+(\mathbf{H}S_{1})^{\prime}\mathbf{I}_{T}\otimes\Omega^{-1}(\mathbf{H}S_{1}),$
$\bm{\mu}_{b}$ $=K_{b}^{-1}\left(\left(\mathbf{I}_{T}\otimes\mathbf{X}(\lambda)\right)^{\prime}\Omega_{1}^{-1}(\mathbf{y}-\theta_{\tau_{1}}\sigma\mathbf{w})+S_{1}\mathbf{H}^{\prime}\mathbf{I}_{T}\otimes\Omega^{-1}\left(\bm{\gamma}-\mathbf{H}S_{2}z\right)\right).$
4 Empirical Application
This section is devoted to studying the effectiveness of the proposed TVP-QR-NS model for zero-coupon US treasury yields. In particular, we split the following section into an in-sample analysis of the yield curve factor and a forecasting exercise.
4.1 Data
Our empirical investigation employs zero-coupon US Treasury yields as constructed by Liu and Wu (2021) through a non-parametric kernel smoothing methodology. The application of this novel yield curve dataset by Liu and Wu (2021) is demonstrated to offer a more precise depiction of the underlying data when compared to the alternative metric presented by Gürkaynak et al. (2007). Following the methodology established by Diebold et al. (2006), our analysis encompasses zero-coupon US Treasury maturities spanning 3, 6, 9, 12, 15, 18, 21, 24, 30, 36, 48, 60, 72, 84, 96, 108, and 120 months. Additionally, our investigation includes key monthly macroeconomic indicators — Industrial Production, CPI Inflation, and the Fed Funds Rate — in line with the framework suggested by Diebold et al. (2006). These macroeconomic variables are chosen to capture the dynamic aspects of the macroeconomy and are all sourced from the US FRED database. Both Industrial Production and CPI Inflation were transformed to log-differenced growth rates. Our study period extends from November 1971 to December 2022.
4.2 In-Sample Results
Previous studies, including Diebold and Li (2006); Diebold et al. (2006, 2008); Mumtaz and Surico (2009); Mönch (2012), focused on the yield curve and its effects within a conditional mean framework. Our study is distinct from the former as it is the first to explore the yield curve from a quantile or distributional perspective. This section presents our empirical in-sample results. The first subsection details the posterior estimates of yield curve factors for each quantile level. The second subsection analyses the distributional features of short- and long-term interest rates during recent crises (the Great Recession and the COVID-19 pandemic). Lastly, the third subsection investigates the interaction between yield curve factors and the macroeconomy from a quantile perspective.
4.2.1 Quantile Yield Curve Factor Estimates
Figures 2, 3, and 4 depict the posterior mean estimates of time-varying yield curve factors at five quantiles (10%, 25%, 50%, 75%, 90%), providing insights into their dynamic evolution. A consistent finding across all graphs is the presence of substantial temporal variation. Beginning with the level factor, interpreted as the long-term interest rate (i.e., yield as maturity approaches infinity), a steady decline in this rate is discernible across all quantiles since the oil price crisis of the 1980s. Figure 2 reveals a distinct pattern: following both the dot-com recession in the early 2000s and the Great Recession of 2008-09, the posterior estimates for the 90th and median quantiles closely align, while those for the 10th quantile diverge. Conversely, after the COVID-19 recession, the opposite trend emerged, with the 10th and median quantiles exhibiting proximity while the 90th quantile diverged. This evidence suggests that post the early 2000s recession and the Great Recession, the distribution of the long-term interest rate skewed leftwards. In contrast, after the COVID-19 recession, the distribution skewed rightwards, likely attributed to the pronounced inflationary pressures experienced in the aftermath of the COVID-19 period.
Figure 2: Posterior mean estimates of the level factor for each of the five quantiles (10%, 25%, 50%, 75%, 90%). NBER recession periods in grey shades.
Turning to the slope and curvature of the yield curve in Figures 3 and 4, the degree of uncertainty associated with these factors varies over time. Notably, the curvature factor exhibits higher volatility across all quantiles than its slope counterpart. [^1] Notice that crossing of the estimated conditional quantiles for the curvature factor is not harmful since it does not have a direct link to one particular yield. A distinct pattern is evident before 1985, where the posterior estimates for the slope factor across all five quantiles are closely clustered, indicating a period of low uncertainty. In contrast, during the oil price crisis of the 1980s, the curvature factor displayed substantial dispersion across quantiles.
Figure 3: Posterior mean estimates of the slope factor for each of the five quantiles (10%, 25%, 50%, 75%, 90%). NBER recession periods in grey shades.
Post-recessionary periods following the early 1990s, the 2000s, and the Great Recession of 2008-09 reveal elevated uncertainty in both slope and curvature factors. Specifically, convergence is observed between the median and the 90th quantile, with the 10th quantile exhibiting a discernible deviation from this convergent trajectory. In the aftermath of the COVID-19 recession, a reverse trajectory emerges, with the 90th quantile diverging from the 10th and median quantiles. The associated uncertainty surrounding both factors appears comparatively subdued compared to the preceding recessions of the 2000s. Both the early 2000s recession and the Great Recession were attributed to financial crises. Our empirical findings suggest a potential connection between periods following financial crises and increased uncertainty in the slope and curvature of the yield curve.
Figure 4: Posterior mean estimates of the curvature factor for each of the five quantiles (10%, 25%, 50%, 75%, 90%). NBER recession periods in grey shades.
It is worthwhile mentioning that these results could not be obtained using a conditional mean approach, which is instead able to produce estimates of the mean yield factors over time. Conversely, the proposed quantile approach enables deeper and more detailed analyses of the term structure while maintaining a low computational cost.
4.2.2 The Distributional Features of the Short- and Long-Term Interest Rate
The proposed novel framework provides a distinct advantage by facilitating the investigation of distributional characteristics in both short- and long-term interest rates over time. As mentioned earlier, the long-term interest rate is identified as the level factor, while the short-term interest rate is a composite of the level and slope factors. We adhere to the methodology outlined by Mitchell et al. (2022) to construct the distributions of these interest rates. Employing an equal weighting scheme, we aggregate all MCMC draws across five quantiles and apply the Epanechnikov kernel, as proposed by Gaglianone and Lima (2012).
The previous analysis revealed an increase in the degree of dispersion among quantiles during post-recessionary periods, particularly in the 2000s. Expanding on this finding, we investigate whether the distributional features of short- and long-term interest rates exhibit temporal variations across the two most recent recessions. Figure 5 illustrates the distributions of both short- and long-term interest rates for selected months preceding and following the Great Recession and the COVID-19 recession.
Figure 5: distribution of the short-term and long-term (level) factors during the Great Recession (left) and the COVID-19 pandemic (right).
Focusing on the short-term interest rate first, discernible evidence of a pronounced distributional shift is observed across the three selected months for both recessions. Notably, during the Great Recession (September 2009), a manifestation of multi-modality in the distribution is apparent. Additionally, after the COVID-19 recession (December 2020), the distribution became less symmetric and more negatively skewed to the left. Turning to the long-term interest rate, the evidence of distribution shifts is more prominent during the COVID-19 recession than during the Great Recession. The long-term interest rate distribution exhibits increased leftward skewness during the Great Recession (September 2009) and the subsequent period (December 2009). Conversely, during the COVID-19 recession, the distribution becomes more positively skewed to the right. Therefore, our results highlight that both short- and long-term interest rate distributions display heightened skewness and asymmetric shapes during recessionary periods. Thus highlighting the importance of using quantile models instead of standard Nelson-Siegel models that are not able to detect the left and right tails of the distribution.
4.2.3 The Interaction between the Yield Curve and Macroeconomy - Quantile VAR Parameter estimates
The proposed framework, as delineated in eq. (9), introduces an avenue for interaction between yield curve factors and macroeconomic variables by incorporating a VAR-like state equation. The estimated off-diagonal VAR coefficients for various quantile levels are presented in Figure 6, shedding light on the directional effects between yield curve factors and macroeconomic variables.
Panel A
Panel B
Figure 6: Posterior mean and credible intervals for the VAR coefficients across different quantile levels.
Panel A of the figure conveys the directional impact of yield curve factors on macroeconomic variables, while Panel B reports the converse scenario. The patterns of influence of each yield factor on the three macroeconomic variables exhibit remarkable similarity. Precisely, evidence of a U-shaped effect across quantiles is discerned for industrial production and inflation, with an inverse relationship observed for the Fed funds rate. Conversely, the impact of macroeconomic variables on the yield curve factor demonstrates higher heterogeneity. Industrial production exerts a significant positive effect on both level and curvature, whereas its impact on the slope is mixed. The effects of inflation exhibit qualitative similarity, albeit with a smaller magnitude on the level and a higher impact on the slope. In contrast, the Fed funds rate positively influences all yield curve factors. Analysing the relationship with quantile levels, the effects on the slope exhibit a U-shaped pattern for industrial production and inflation, while the Fed funds rate demonstrates an inverse U-shaped relationship. Meanwhile, the relationship with the level remains constant across quantiles. In summary, our findings provide substantial evidence that the interaction between yield curve factors and the macroeconomy significantly varies across quantiles.
4.3 Forecasting Yields Tail Risk
We conduct a pseudo-out-of-sample forecasting exercise to assess the utility of our novel quantile Nelson-Siegel approach, comparing it against the corresponding standard Nelson-Siegel conditional mean counterpart for forecasting tail risks in zero-coupon yields. The initial holdout period spans from November 1971 to January 2000, with a subsequent forecast evaluation period extending from February 2000 to December 2022 (comprising 264 evaluation periods). We generate 1st, 6th, and 12th step-ahead tail risk forecasts and employ the quantile score (QS) (see Giacomini and Komunjer, 2005; Carriero et al., 2022) as the forecast evaluation metric to gauge the performance of the two models. Following Gneiting and Ranjan (2011), we define the QS as:
$QS_{\tau,i,t}=(y_{t}(m)-\mathcal{Q}_{\tau,i,t})-(\tau-\mathbb{I}\{y_{t}(m)\leq\mathcal{Q}_{\tau,i,t}\}).$ (17)
Here, $QS_{\tau,i,t}$ denotes the predictive quantile of the $m$-th yield, and $\mathbb{I}\{y_{t}(m)\leq\mathcal{Q}_{\tau,i,t}\}$ takes the value 1 if the realised value is at or below the predictive quantile and 0 otherwise. We evaluate the QS in the upper and lower tails by setting $\tau=0.9$ and $\tau=0.1$, respectively.
Table 1 presents the 10th and 90th percentiles of QS for the out-of-sample forecasting exercise. The top panel of Table 1 displays the average QS of our proposed approach relative to their corresponding conditional mean counterparts for each of the 17 yields across the evaluation period. A quantile score less than 1 implies that our proposed approach outperforms the corresponding conditional mean counterpart. Notably, our proposed approach consistently produces superior tail forecasts for all 17 yields and across all forecast horizons compared to their corresponding conditional mean counterparts. We further assess the significance of our forecasting results through a Diebold-Mariano test, with results presented in the bottom panel of Table 1, clearly indicating the statistical significance of our forecasting results in favour of our proposed approach.
Additionally, we depict the average raw QS across the 17 yields for both models over time in Figure 7. The top and bottom panels show the 10th and 90th percentiles of QS, respectively. Three key features emerge from this figure. First, reinforcing the results in Table 1, our proposed quantile Nelson-Siegel model consistently outperforms their corresponding conditional mean counterparts across all horizons and quantile levels, with the evidence being stronger for 1-step ahead forecasts. Second, the 1-step ahead QS (for both 10th and 90th percentiles) for the conditional mean model exhibits significantly more instability over time than our quantile approach. Finally, we find strong evidence in favour of the proposed time-varying parameters quantile Nelson-Siegel model for forecasting the left tail of the yield distribution, suggesting its ability to more appropriately capture the lower bound of the yield curve compared to standard conventional models.
Figure 7: Quantile Score over time for the 1-, 6-, and 12-steps ahead forecast horizons, at quantile levels 10% (top plots) and 90% (bottom plots), for the QR-NS model (blue) and the standard NS model (red). Shaded areas represent NBER recession periods.
Table 1: Forecasting results for the quantile NS relative to standard NS model, together with Diebold-Mariano test. Out-of-sample evaluation period: 2000 - 2022.
Ratio of quantile scores (quantile NS over standard NS)
10% QS
Horizon 3 m 6 m 9 m 12 m 15 m 18 m 21 m 24 m 30 m 36 m 48 m 60 m 72 m 84 m 96 m 108 m 120 m Average
1-step ahead 0.28 0.34 0.39 0.43 0.44 0.45 0.45 0.45 0.43 0.42 0.41 0.42 0.45 0.50 0.54 0.60 0.65 0.45
6-step ahead 0.61 0.63 0.64 0.65 0.65 0.65 0.64 0.64 0.63 0.61 0.59 0.58 0.57 0.57 0.57 0.58 0.59 0.61
12-step ahead 0.67 0.68 0.69 0.70 0.70 0.71 0.71 0.70 0.70 0.69 0.67 0.66 0.65 0.63 0.63 0.62 0.62 0.67
90% QS
Horizon 3 m 6 m 9 m 12 m 15 m 18 m 21 m 24 m 30 m 36 m 48 m 60 m 72 m 84 m 96 m 108 m 120 m Average
1-step ahead 0.47 0.55 0.56 0.55 0.53 0.51 0.50 0.49 0.49 0.49 0.50 0.53 0.54 0.56 0.57 0.57 0.56 0.53
6-step ahead 0.98 0.98 0.96 0.93 0.90 0.88 0.86 0.84 0.80 0.77 0.73 0.71 0.69 0.69 0.69 0.69 0.68 0.81
12-step ahead 0.96 0.96 0.95 0.93 0.92 0.91 0.89 0.88 0.86 0.84 0.82 0.80 0.79 0.78 0.77 0.76 0.75 0.86
Diebold-Mariano test, significance level: 1 (10%), 2 (5%), 3 (1%)
10% QS
Horizon 3 m 6 m 9 m 12 m 15 m 18 m 21 m 24 m 30 m 36 m 48 m 60 m 72 m 84 m 96 m 108 m 120 m Average
1-step ahead 3.00 3.00 3.00 3.00 3.00 3.00 3.00 3.00 3.00 3.00 3.00 3.00 3.00 3.00 3.00 3.00 3.00 3.00
6-step ahead 3.00 3.00 3.00 3.00 3.00 3.00 3.00 3.00 3.00 3.00 3.00 3.00 3.00 3.00 3.00 3.00 3.00 3.00
12-step ahead 3.00 3.00 3.00 3.00 3.00 3.00 3.00 3.00 3.00 3.00 3.00 3.00 3.00 3.00 3.00 3.00 3.00 3.00
90% QS
Horizon 3 m 6 m 9 m 12 m 15 m 18 m 21 m 24 m 30 m 36 m 48 m 60 m 72 m 84 m 96 m 108 m 120 m Average
1-step ahead 3.00 3.00 3.00 3.00 3.00 3.00 3.00 3.00 3.00 3.00 3.00 3.00 3.00 3.00 3.00 3.00 3.00 3.00
6-step ahead 0.00 0.00 2.00 3.00 3.00 3.00 3.00 3.00 3.00 3.00 3.00 3.00 3.00 3.00 3.00 3.00 3.00 2.59
12-step ahead 1.00 1.00 2.00 3.00 3.00 3.00 3.00 3.00 3.00 3.00 3.00 3.00 3.00 3.00 3.00 3.00 3.00 2.71
5 Conclusions
In this study, we introduce a novel time-varying parameters quantile Nelson-Siegel (TVP-QR-NS) model within a Bayesian framework and apply it to an empirical analysis of the US term structure of interest rates. Our in-sample findings reveal an increased dispersion in yield curve factors across quantiles in the post-recession period. Additionally, temporal variations in the distribution of short- and long-term interest rates are observed during the two most recent recessions. Furthermore, evidence indicates significant fluctuations in the interaction between yield curve factors and the macroeconomy across quantiles. Finally, through an out-of-sample forecasting exercise, we demonstrate the superior performance of our proposed approach in tail risk forecasting across all yields compared to its standard conditional mean counterpart.

### v2

@@
-A Quantile Nelson-Siegel model
+
 Abstract
-A widespread approach to modelling the interaction between macroeconomic variables and the yield curve relies on three latent factors usually interpreted as the level, slope, and curvature (Diebold et al., 2006).
-This approach is inherently focused on the conditional mean of the yields and postulates a dynamic linear model where the latent factors smoothly change over time.
-However, periods of deep crisis, such as the Great Recession and the recent pandemic, have highlighted the importance of statistical models that account for asymmetric shocks and are able to forecast the tails of a variable’s distribution.
-A new version of the dynamic three-factor model is proposed to address this issue based on quantile regressions.
-The novel approach leverages the potential of quantile regression to model the entire (conditional) distribution of the yields instead of restricting to its mean.
-An application to US data from the 1970s shows the significant heterogeneity of the interactions between financial and macroeconomic variables across different quantiles.
-Moreover, an out-of-sample forecasting exercise showcases the proposed method’s advantages in predicting the yield distribution tails compared to the standard conditional mean model.
-Finally, by inspecting the posterior distribution of the three factors during the recent major crises, new evidence is found that supports the greater and longer-lasting negative impact of the great recession on the yields compared to the COVID-19 pandemic.
-Keywords: Bayesian inference; Nelson-Siegel model; Quantile regression; Time-varying parameters; Yield curve.
+We propose a novel framework for modeling the yield curve from a quantile perspective.
+Building on the dynamic Nelson–Siegel model of Diebold et al.
+(2006), we extend its traditional mean-based approach to a quantile regression setting, enabling the estimation of yield curve factors—level, slope, and curvature—at specific quantiles of the conditional distribution.
+A key advantage of our framework is its ability to characterize the entire conditional distribution of the yield curve across maturities and over time.
+In an empirical analysis of the U.S.
+term structure of interest rates, our method demonstrates superior out-of-sample forecasting performance, particularly in capturing the tails of the yield distribution—an aspect increasingly emphasized in the recent literature on distributional forecasting.
+In addition to its forecasting advantages, our approach reveals rich distributional features beyond the mean.
+In particular, we find that the dynamic changes in these distributional features differ markedly between the Great Recession and the COVID-19 pandemic period, highlighting a fundamental shift in how interest rate markets respond to distinct economic shocks.
+Keywords: Nelson-Siegel; Yield Curve; Bayesian Markov Chain Monte-Carlo; Quantile regression.
 1 Introduction
@@
 This feature is particularly crucial when undertaking the modelling and forecasting of interest rate term structures, as emphasised in our empirical application.
-Figure 1: Distributional term structure of the interest rate for June and December 2022.
+Figure 1: Distributional term structure of the interest rate for June (left) and December (right) 2022.
 Inference is performed within a Bayesian framework by leveraging the connection between a likelihood formed by the asymmetric Laplace distribution (Kotz et al., 2001) and quantile regression provided by Yu and Moyeed (2001), and Petrella and Raponi (2019).
@@
 Overall, this combination provides a simple framework and computationally efficient method to investigate and forecast the term structure at every quantile level of interest.
-The proposed method is then applied to investigate and forecast the US term structure of interest rates.
-Specifically, an in-sample analysis highlights the advantages of the quantile approach compared to the standard conditional mean Nelson-Siegel model.
-In particular, we are able to obtain the entire distribution of the three factors at different points in time, thus enabling a more detailed analysis of the effects of the recent crises (the Great Recession and the COVID-19 pandemic) on the term structure.
-Interestingly, these results, which could not be obtained using a conditional mean approach, reveal substantial differences between the impacts of the financial crisis and COVID-19.
-The former shifted leftwards the distribution of the short-term interest rate without meaningful impact on the long-term.
-Conversely, the pandemic had an initial significant adverse effect on both the short- and long-term rates, but the distribution in 2022 shifted rightwards compared to the pre-COVID one.
-The advantages of the proposed TVP-QR-NS model are better highlighted in terms of tail risk yield forecasting.
-Compared to the conditional mean model, our approach provides better out-of-sample forecasts in both tails of the yield distribution, as measured by the quantile score.
-The difference is particularly striking on the left tail and consistent over time.
+Moreover, our novel econometric framework shares similarities with methodologies put forth by Korobilis and Schröder (2024, 2025).
+In the former, the authors propose a variational Bayes approach to estimate a quantile static factor model, while the latter extends the factor-augmented vector autoregression to a quantile framework.
+The primary distinction between our study and both of these lies in our specific focus on analysing the distribution of the yield curve.
+In contrast, their emphasis is on examining macroeconomic uncertainty within the US and macroeconomic risk across multiple countries.
+The proposed method is applied to investigate and forecast the US term structure of interest rates, with particular emphasis on capturing the full distributional dynamics of key yield curve factors.
+To validate its utility, we conduct an out-of-sample forecasting exercise to evaluate the performance of the TVP-QR-NS model against standard conditional mean benchmarks widely used in the empirical literature.
+This study contributes to a growing body of research that focuses on forecasting macroeconomic variables from a distributional perspective (see Carriero et al.
+(2022); Clark et al.
+(2024); Lenza et al.
+(2025)).
+The results indicate that the TVP-QR-NS model consistently delivers superior tail risk forecasts relative to conventional benchmarks, particularly at longer forecast horizons and for long-term maturities.
+Moreover, employing the model confidence set methodology proposed by Hansen et al.
+(2011), we find that the TVP-QR-NS model is systematically retained within the confidence sets for longer maturities and long-term forecast horizons, while standard conditional mean models are frequently excluded, underscoring the robustness and reliability of our approach.
+The in-sample analysis highlights the advantages of the quantile regression framework over the conventional conditional mean Nelson-Siegel model.
+Unlike mean-based frameworks, our method enables estimation of the entire distribution of the three Nelson-Siegel factors, as well as yields at selected maturities over time.
+This comprehensive characterization allows for a more detailed assessment of the effects of recent economic crises, specifically the Great Recession and the COVID-19 pandemic, on the term structure.
+The results uncover critical insights that are not accessible through conditional mean models.
+Specifically, the impact of the Great Recession and COVID-19 on the distribution of short-term interest rates diverges substantially, with clear distributional shifts evident across the two crises.
+In contrast, the distribution of long-term rates during the Great Recession exhibited no significant shift in location, although an increase in left-skewness was observed.
+Further analysis shows that, following the Great Recession, a pronounced location shift occurred across the short, medium, and long ends of the yield curve, with the 10-year maturity exhibiting increased left-skewness.
+Conversely, during the COVID-19 recession, the location shift was confined to the long end of the yield curve and was accompanied by heightened dispersion.
+Moreover, we find robust evidence of quantile-dependent dynamics in the relationship between yield curve factors and macroeconomic conditions, underscoring the importance of capturing heterogeneity across the distribution.
 The remainder of this article is organised as follows.
@@
 for each quantile level $\tau\in(0,1)$ as follows
-$Q_{t}^{\tau}(m)=\beta_{1t}(\tau)+\beta_{2t}(\tau)\Big{(}\frac{1-e^{-\lambda^{\tau}m}}{\lambda^{\tau}m}\Big{)}+\beta_{3t}(\tau)\Big{(}\frac{1-e^{-\lambda^{\tau}m}}{\lambda^{\tau}m}-e^{-\lambda^{\tau}m}\Big{)},$ (3)
-where $m$ is the maturity, $\beta_{t1}^{\tau}$, $\beta_{t2}^{\tau}$, and $\beta_{t3}^{\tau}$ denote the (unobserved) level, slope and curvature factors, and $\lambda^{\tau}\in\mathds{R}_{+}$ is a decay parameter.
-In the following, we shall assume $\lambda^{\tau}=\lambda$.
+$Q_{t}^{\tau}(m)=\beta_{1t}(\tau)+\beta_{2t}(\tau)\left(\frac{1-e^{-\lambda(\tau)m}}{\lambda(\tau)m}\right)+\beta_{3t}(\tau)\left(\frac{1-e^{-\lambda(\tau)m}}{\lambda(\tau)m}-e^{-\lambda(\tau)m}\right),$ (3)
+where $m$ is the maturity, $\beta_{t1}^{\tau}$, $\beta_{t2}^{\tau}$, and $\beta_{t3}^{\tau}$ denote the (unobserved) level, slope and curvature factors, and $\lambda(\tau)\in\mathds{R}_{+}$ is a decay parameter.
 Thus, by directly modelling the quantile function, we allow the yield curve factors to behave differently across the quantile levels.
 Consider a series of observations at specific maturities $m\in\{1,\ldots,M\}$ and denote with $\mathbf{y}_{t}=(y_{t}(m_{1}),\ldots,y_{t}(m_{M}))^{\prime}$ the vector of yields at different maturity.
-Let $\mathbf{x}(m;\lambda)=\big{(}1,\frac{1-e^{-\lambda m}}{\lambda m},\frac{1-e^{-\lambda m}}{\lambda m}-e^{-\lambda m}\big{)}^{\prime}$, and $\mathbf{X}(\lambda)=\left[\mathbf{x}(m_{1};\lambda),\mathbf{x}(m_{2};\lambda),\ldots,\mathbf{x}(m_{M};\lambda)\right]^{\prime}\in\mathds{R}^{M\times 3}$.
+Let $\mathbf{x}(m;\lambda(\tau))=\big{(}1,\frac{1-e^{-\lambda(\tau)m}}{\lambda(\tau)m},\frac{1-e^{-\lambda(\tau)m}}{\lambda(\tau)m}-e^{-\lambda(\tau)m}\big{)}^{\prime}$, and $,\mathbf{X}(\lambda(\tau))=[\mathbf{x}(m_{1};\lambda(\tau)),\mathbf{x}(m_{2};\lambda(\tau)),$$\ldots,\mathbf{x}(m_{M};\lambda(\tau))]^{\prime}\in\mathds{R}^{M\times 3}$.
 Then, the conditional quantile model can be written as
-$Q^{\tau}_{t}=\mathbf{X}(\lambda)\bm{\beta}_{t}(\tau)$ (4)
+$Q^{\tau}_{t}=\mathbf{X}(\lambda(\tau))\bm{\beta}_{t}(\tau)$ (4)
 such that $Q^{\tau}_{t}=(Q^{\tau}_{t}(m_{1}),\ldots,Q^{\tau}_{t}(m_{M}))^{\prime}$ and $\bm{\beta}_{t}(\tau)=(\beta_{1t}(\tau),\beta_{2t}(\tau),\beta_{3t}(\tau))^{\prime}$.
@@
 Differently from Coroneo et al.
-(2016) that proposed an approach related to ours in the context of a conditional mean model for the yields, we are modelling the conditional quantile of $y_{t}(m)$ and rely on an alternative specification of the interactions between the yield and the macroeconomic factors.
+(2016) that proposed an approach related to our in the context of a conditional mean model for the yields, we are modelling the conditional quantile of $y_{t}(m)$ and rely on an alternative specification of the interactions between the yield and the macroeconomic factors.
 To design a regression model for the yield such that the conditional $\tau$-quantile corresponds to eq.
@@
 where $\rho_{\tau}(x)=x(\tau-\mathbb{I}(x\leq\tau))$ is the check-loss function. Therefore, combining the conditional quantile in eq. (3) with the asymmetric Laplace innovation results in
-$y_{t}(m)=\mathbf{x}(m;\lambda)^{\prime}\bm{\beta}_{t}(\tau)+\epsilon_{t}(m),\qquad\epsilon_{t}(m)\sim\mathcal{AL}(0,\sigma,\tau).$ (6)
+$y_{t}(m)=\mathbf{x}(m;\lambda(\tau))^{\prime}\bm{\beta}_{t}(\tau)+\epsilon_{t}(m),\qquad\epsilon_{t}(m)\sim\mathcal{AL}(0,\sigma,\tau).$ (6)
 It is important to highlight that the asymmetric Laplace distribution admits a stochastic representation as a location-scale mixture of Gaussian (Kotz et al., 2001), thus allowing to rewrite eq.
 (6) as
-$y_{t}(m)\stackrel{{\scriptstyle d}}{{=}}\mathbf{x}(m;\lambda)^{\prime}\bm{\beta}(\tau)+w_{m,t}\sigma\theta_{\tau,1}+\sqrt{w_{m,t}}\sigma\theta_{\tau,2}u_{t}(m),\qquad u_{t}(m)\sim\mathcal{N}(0,1),$ (7)
-where $w_{m,t}\sim\mathcal{E}xp(1)$ is exponential with rate one.
+$y_{t}(m)\stackrel{{\scriptstyle d}}{{=}}\mathbf{x}(m;\lambda(\tau))^{\prime}\bm{\beta}_{t}(\tau)+w_{t}(m)\sigma\theta_{\tau,1}+\sqrt{w_{t}(m)}\sigma\theta_{\tau,2}u_{t}(m),\quad u_{t}(m)\sim\mathcal{N}(0,1),$ (7)
+where $w_{t}(m)\sim\mathcal{E}xp(1)$ is exponential with rate one.
 To establish an equivalence between the mode of the asymmetric Laplace distribution and the optimisation of the standard quantile regression problem in Koenker and Bassett (1978), the parameters $\theta_{\tau,1},\theta_{\tau,2}$ are constrained to the following values (Kotz et al., 2001):
 $\theta_{\tau,1}=\frac{1-2\tau}{\tau(1-\tau)},\qquad\theta_{\tau,2}=\sqrt{\frac{2}{\tau(1-\tau)}}.$ (8)
-Throughout the article, we follow the existing literature and fix the decay parameter at $\lambda=0.0609$ (e.g., see Diebold and Li, 2006).
-Moreover, the scale of the innovation is set to $\sigma=1.0$, as in the original quantile regression model (Yu and Moyeed, 2001).
+In the existing literature, the decay parameter is often fixed at $\lambda=0.0609$ (e.g., Diebold and Li, 2006).
+In contrast, we adopt an agnostic approach by estimating $\lambda(\tau)$ separately for each quantile, and assume a uniform prior distribution, $\lambda(\tau)\sim\mathcal{U}(0,0.3)$ (see the appendix for further details).
+Given that the conditional posterior distribution of $\lambda(\tau)$ is non-standard, we employ a Griddy-Gibbs sampling step to draw from its posterior.
+Finally, the scale of the innovation is set to $\sigma=1.0$, as in the original quantile regression model (Yu and Moyeed, 2001).
 The combination of eq. (5) and (6) provides the state-space form of the proposed time-varying parameters quantile Nelson-Siegel model (TVP-QR-NS) as
-$\begin{split}y_{t}(m)&=\mathbf{x}(m;\lambda)^{\prime}\bm{\beta}_{t}(\tau)+\epsilon_{t}(m),\qquad\epsilon_{t}(m)\sim\mathcal{AL}(0,\sigma,\tau)\\ \bm{\phi}_{t}&=\bm{\gamma}_{0}+\Gamma\bm{\phi}_{t-1}+\bm{\eta}_{t},\qquad\bm{\eta}_{t}\sim\mathcal{N}(\mathbf{0},\Omega),\end{split}$ (9)
-where we recall the fixed values $\lambda=0.0609$ and $\sigma=1.0$.
+$\begin{split}y_{t}(m)&=\mathbf{x}(m;\lambda(\tau))^{\prime}\bm{\beta}_{t}(\tau)+\epsilon_{t}(m),\qquad\epsilon_{t}(m)\sim\mathcal{AL}(0,\sigma,\tau)\\ \bm{\phi}_{t}&=\bm{\gamma}_{0}+\Gamma\bm{\phi}_{t-1}+\bm{\eta}_{t},\qquad\bm{\eta}_{t}\sim\mathcal{N}(\mathbf{0},\Omega),\end{split}$ (9)
 Denoting with $\mathbf{Y}=(\mathbf{y}_{1},\ldots,\mathbf{y}_{T})^{\prime}\in\mathds{R}^{T\times M}$ the matrix of observations, the joint density of the observables and auxiliary variables (conditional on the time-varying factors) is
-$p(\mathbf{Y},\mathbf{w}|\bm{\beta})$ $=\prod_{m=1}^{M}\mathcal{E}xp(w_{m,t}|1)\times\mathcal{N}\big{(}y_{t}(m)|\mathbf{x}(m;\lambda)^{\prime}\bm{\beta}_{t}(\tau)+w_{m,t}\sigma\theta_{\tau,1},\>w_{m,t}\sigma^{2}\theta_{\tau,2}^{2}\big{)}.$ (10)
+$p(\mathbf{Y},\mathbf{w}|\bm{\beta})$ $=\prod_{m=1}^{M}\mathcal{E}xp(w_{t}(m)|1)\times\mathcal{N}\big{(}y_{t}(m)|\mathbf{x}(m;\lambda(\tau))^{\prime}\bm{\beta}_{t}(\tau)+w_{t}(m)\sigma\theta_{\tau,1},\>w_{t}(m)\sigma^{2}\theta_{\tau,2}^{2}\big{)}.$ (10)
 3 Estimation
@@
 By leveraging the mixture representation of the asymmetric Laplace distribution, we obtain a conditionally linear Gaussian state space for the latent dynamic vector $\bm{\beta}_{t}$.
-Thus, we sample from the posterior full conditional distribution of $\bm{\beta}_{t}$ directly in a single loop, using a precision sampler (e.g., see Chan and Jeliazkov, 2009; Chan et al., 2020) as described in the following subsection.
-Let us define $e_{m,t}=y_{t}(m)-\mathbf{x}(m;\lambda)^{\prime}\bm{\beta}_{t}$. Then, the posterior full conditional distribution of the auxiliary variable $w_{t}$ is obtained as
-$w_{m,t}|y_{t}(m),\bm{\beta}_{t}$ $\propto w_{m,t}^{-\frac{1}{2}}\exp\Big{\{}-\frac{1}{2}\Big{[}\frac{\theta_{\tau,1}^{2}}{\theta_{\tau,2}^{2}}w_{m,t}+2w_{m,t}+w_{m,t}^{-1}\frac{e_{m,t}^{2}}{\theta_{\tau,2}^{2}\sigma^{2}}\Big{]}\Big{\}},$ (12)
+Thus, we sample the full conditional posterior distribution of $\bm{\beta}_{t}$ directly in a single loop, using a precision sampler (e.g., see Chan and Jeliazkov, 2009; Chan et al., 2020) as described in the following subsection.
+Let us define $e_{t}(m)=y_{t}(m)-\mathbf{x}(m;\lambda(\tau))^{\prime}\bm{\beta}_{t}$. Then, the full conditional posterior distribution of the auxiliary variable $w_{t}(m)$ is obtained as
+$w_{t}(m)|y_{t}(m),\bm{\beta}_{t}$ $\propto w_{t}(m)^{-\frac{1}{2}}\exp\left\{-\frac{1}{2}\left[\frac{\theta_{\tau,1}^{2}}{\theta_{\tau,2}^{2}}w_{t}(m)+2w_{t}(m)+w_{t}(m)^{-1}\frac{e_{t}(m)^{2}}{\theta_{\tau,2}^{2}\sigma^{2}}\right]\right\},$ (12)
 which is the kernel of the generalized inverse Gaussian distribution $\text{GiG}(\overline{p}_{w},\overline{a}_{w},\overline{b}_{w,m,t})$ with
-$\overline{p}_{w}=1-\frac{1}{2},\qquad\overline{a}_{w}$ $=2+\frac{\theta_{\tau,1}^{2}}{\theta_{\tau,2}^{2}},\qquad\overline{b}_{w,m,t}=\frac{e_{m,t}^{2}}{\theta_{\tau,2}^{2}\sigma^{2}}.$
+$\overline{p}_{w}=1-\frac{1}{2},\qquad\overline{a}_{w}$ $=2+\frac{\theta_{\tau,1}^{2}}{\theta_{\tau,2}^{2}},\qquad\overline{b}_{w,m,t}=\frac{e_{t}(m)^{2}}{\theta_{\tau,2}^{2}\sigma^{2}}.$
 Finally, we specify a Gaussian prior for the vectorised coefficient matrix of the state equation, $\bm{\gamma}=(\bm{\gamma}_{0}^{\prime},\operatorname{vec}(\Gamma)^{\prime})^{\prime}$, and an inverse Wishart prior for the innovation covariance, that is
@@
 Specifically, we define the stacked vectors $\mathbf{y}=(\mathbf{y}_{1}^{\prime},\ldots,\mathbf{y}_{T}^{\prime})^{\prime}$, $\bm{\beta}=(\bm{\beta}_{1}(\tau)^{\prime},\ldots,\bm{\beta}_{T}(\tau)^{\prime})^{\prime}$, $\mathbf{w}=(w_{1},\ldots,w_{T})^{\prime}$, and $\mathbf{u}=(u_{1},\ldots,u_{T})^{\prime}$, we have
-$\mathbf{y}=\mathbf{I}_{T}\otimes\mathbf{X}(\lambda)\bm{\beta}+\theta_{\tau,1}\sigma\mathbf{w}+\mathbf{u},$ (15)
+$\mathbf{y}=\left(\mathbf{I}_{T}\otimes\mathbf{X}(\lambda(\tau))\right)\bm{\beta}+\theta_{\tau,1}\sigma\mathbf{w}+\mathbf{u},$ (15)
 where $\mathbf{u}\sim\mathcal{N}(\mathbf{0}_{TM},\sigma^{2}\theta_{\tau,2}\Omega_{1})$ with $\Omega_{1}=\operatorname{diag}(\mathbf{w})$.
@@
 Thus, we obtain the conditional posterior of the time-varying yield curve factors as $\bm{\beta}|\mathbf{Y},\mathbf{w},\Gamma,\bm{\gamma}_{0},\Omega\propto\mathcal{N}\left(\bm{\mu}_{b},K_{b}^{-1}\right)$, where
-$K_{b}$ $=\frac{1}{\sigma^{2}\theta_{\tau,2}}\left(\mathbf{I}_{T}\otimes\mathbf{X}(\lambda)\right)^{\prime}\Omega_{1}^{-1}\left(\mathbf{I}_{T}\otimes\mathbf{X}(\lambda)\right)+(\mathbf{H}S_{1})^{\prime}\mathbf{I}_{T}\otimes\Omega^{-1}(\mathbf{H}S_{1}),$
-$\bm{\mu}_{b}$ $=K_{b}^{-1}\left(\left(\mathbf{I}_{T}\otimes\mathbf{X}(\lambda)\right)^{\prime}\Omega_{1}^{-1}(\mathbf{y}-\theta_{\tau_{1}}\sigma\mathbf{w})+S_{1}\mathbf{H}^{\prime}\mathbf{I}_{T}\otimes\Omega^{-1}\left(\bm{\gamma}-\mathbf{H}S_{2}z\right)\right).$
+$K_{b}$ $=\frac{1}{\sigma^{2}\theta_{\tau,2}}\big{(}\mathbf{I}_{T}\otimes\mathbf{X}(\lambda(\tau))\big{)}^{\prime}\Omega_{1}^{-1}\left(\mathbf{I}_{T}\otimes\mathbf{X}(\lambda)\right)+S_{1}\mathbf{H}^{\prime}\big{(}\mathbf{I}_{T}\otimes\Omega^{-1}\big{)}\mathbf{H}S_{1},$
+$\bm{\mu}_{b}$ $=K_{b}^{-1}\Big{(}\big{(}\mathbf{I}_{T}\otimes\mathbf{X}(\lambda(\tau))\big{)}^{\prime}\Omega_{1}^{-1}(\mathbf{y}-\theta_{\tau_{1}}\sigma\mathbf{w})+S_{1}\mathbf{H}^{\prime}\big{(}\mathbf{I}_{T}\otimes\Omega^{-1}\big{)}\left(\bm{\gamma}-\mathbf{H}S_{2}z\right)\Big{)}.$
+3.3 Monotonicity
+Traditional estimation of quantile regression, including our strategy, involves fitting the quantile NS model separately at each quantile level, treating each quantile as an independent modeling task.
+Although this approach provides flexibility and ease of implementation, it overlooks the intrinsic monotonicity that should characterize conditional quantile functions by construction.
+As a result, the estimated quantile curves may violate the logical ordering across quantiles, potentially leading to internal inconsistencies and undermining the interpretability of the distributional forecasts.
+To ensure monotonicity across quantile levels $\tau\in\left\{\tau_{1},...,\tau_{K}\right\}$, it is essential to move beyond separate estimations and instead estimate the quantile NS models jointly across $\tau$.
+This joint estimation framework requires imposing explicit monotonicity constraints across quantiles,
+$\mathbf{x}\big{(}m;\lambda(\tau_{k})\big{)}^{\prime}\bm{\beta}_{t}(\tau_{k})\leq\mathbf{x}\big{(}m;\lambda(\tau_{k+1})\big{)}^{\prime}\bm{\beta}_{t}(\tau_{k+1})$
+for $m=1,\ldots,M,k=1,\ldots,K$ and $t=1,\ldots T$, thus preserving the fundamental ordering structure inherent to conditional quantile functions.
+However, this approach introduces a large number of cross-quantile constraints, and the resulting optimization problem becomes highly complex and computationally burdensome, to the point where joint estimation under monotonicity constraints can be practically infeasible in its full form.
+This poses a major challenge for implementing a theoretically coherent yet computationally tractable distributional forecasting framework.
+To construct a coherent predictive density from the individually estimated quantile NS models, we adopt the re-ordering technique proposed by Chernozhukov et al.
+(2010), a standard approach in the literature for addressing the issue of quantile crossing.
+In the context of the NS model, the yield curve is endowed with additional structural interpretation that can inform both estimation and inference.
+Specifically, the level factor in the NS model captures the long-run component of interest rates and asymptotically corresponds to the yield at extremely long maturities—essentially representing the ultimate long-term rate.
+Conversely, the sum of the level and slope factors determines the instantaneous (short-term) interest rate, reflecting the yield at the shortest maturity.
+Based on these insights, we can potentially impose the following monotonicity constraints
+$\bm{\beta}_{t,1}(\tau_{k})\leq\bm{\beta}_{t,1}(\tau_{k+1})$ (17)
+and
+$\bm{\beta}_{t,1}(\tau_{k})+\bm{\beta}_{t,2}(\tau_{k})\leq\bm{\beta}_{t,1}(\tau_{k+1})+\bm{\beta}_{t,2}(\tau_{k+1}).$ (18)
+Instead of directly imposing these monotonicity constraints within a joint estimation framework—such as through constrained MCMC, which would be computationally intensive—we adopt a more practical post-processing strategy.
+Specifically, we formulate a quadratic programming problem that adjusts the unconstrained estimates of $\bm{\beta}_{t,1}(\tau_{k})$ and $\bm{\beta}_{t,2}(\tau_{k})$ across quantile levels.
+The objective is to find new values that remain as close as possible to the originally estimated posterior mean, $\widehat{\bm{\beta}}_{t,1}$ and $\widehat{\bm{\beta}}_{t,2}$, while satisfying the monotonicity constraints.
+This post-estimation correction ensures internal consistency and preserves the structural interpretation of the NS model, while avoiding the computational burden of imposing constraints during the sampling phase.
 4 Empirical Application
-This section is devoted to studying the effectiveness of the proposed TVP-QR-NS model for zero-coupon US treasury yields.
-In particular, we split the following section into an in-sample analysis of the yield curve factor and a forecasting exercise.
+This section focuses on evaluating the performance of the proposed TVP-QR-NS model applied to zero-coupon US Treasury yields.
+Specifically, the analysis is divided into two parts: an out-of-sample forecasting exercise and an in-sample assessment.
 4.1 Data
@@
 Both Industrial Production and CPI Inflation were transformed to log-differenced growth rates.
-Our study period extends from November 1971 to December 2022.
+Our study period extends from November 1971 to December 2024.
+4.2 Forecasting Yields Tail Risk
+To validate the utility of our proposed framework, we conduct a pseudo-out-of-sample forecasting exercise to evaluate the performance of the quantile Nelson-Siegel approach in capturing tail risks in zero-coupon yields.
+The benchmark model employed for comparison is a univariate autoregressive model with one lag.
+Our proposed model’s performance is assessed relative to this baseline as well as three alternative conditional mean models, as detailed in Table 1.
+These competing specifications include: (i) a time-varying parameter model with stochastic volatility (SV); (ii) the large Bayesian vector autoregression (VAR) with common stochastic volatility (CSV) developed by Chan (2020); and (iii) the conditional mean formulation of the dynamic Nelson-Siegel model of Diebold et al.
+(2006).
+Table 1: Overview of Competing Models
+Model Description
+AR(1) Autoregressive model with one lag (Benchmark model).
+TVP-AR(1)-SV Time-varying parameter autoregressive model with SV and one lag.
+BVAR(1)-CSV Large Bayesian VAR with one lag and common stochastic volatility.
+DNS Dynamic Nelson-Siegel model of Diebold et al. (2006).
+TVP-QNS Proposed Quantile-based time-varying Nelson-Siegel model.
+The initial holdout period spans from November 1971 to January 2000, with a subsequent forecast evaluation period extending from February 2000 to December 2024.
+We generate 1st, 12th, and 36th month-ahead tail risk forecasts and evaluate the predictive performance of the competing models using the quantile score (QS) (see Giacomini and Komunjer, 2005; Carriero et al., 2022).
+Following Gneiting and Ranjan (2011), we define the QS as:
+$QS_{\tau,i,t}=(y_{t}(m)-\mathcal{Q}_{\tau,i,t})-(\tau-\mathbb{I}\{y_{t}(m)\leq\mathcal{Q}_{\tau,i,t}\}).$ (19)
+Here, $\mathcal{Q}_{\tau,i,t}$ denotes the predictive quantile of the $m$-th yield, and $\mathbb{I}\{y_{t}(m)\leq\mathcal{Q}_{\tau,i,t}\}$ takes the value 1 if the realised value is at or below the predictive quantile and 0 otherwise.
+We evaluate the QS in the upper and lower tails by setting $\tau=0.9$ and $\tau=0.1$, respectively.
+Table 2: Average 10% (Lower-Tail) Quantile Scores for Selected Maturities Relative to an AR(1) Benchmark
+Model 3m $p_{\text{MCS}}$ 6m $p_{\text{MCS}}$ 12m $p_{\text{MCS}}$ 24m $p_{\text{MCS}}$ 36m $p_{\text{MCS}}$ 60m $p_{\text{MCS}}$ 120m $p_{\text{MCS}}$ Avg.
+1 Month-Ahead Forecast
+TVP-AR(1)-SV 0.43 0.00 0.42 0.01 0.50 0.03 0.63 0.02 0.72 0.00 0.81 0.00 0.90 0.00 0.68
+BVAR(1)-CSV 0.28 1.00** 0.31 1.00** 0.38 1.00** 0.50 1.00** 0.57 0.64** 0.63 0.23* 0.65 1.00** 0.52
+DNS 1.72 0.00 1.32 0.00 1.06 0.00 1.13 0.00 1.23 0.00 1.29 0.00 1.07 0.00 1.20
+TVP-QNS 0.48 0.00 0.46 0.00 0.47 0.04 0.53 0.46** 0.55 1.00** 0.57 1.00** 0.70 0.41** 0.55
+12 Month-Ahead Forecast
+TVP-AR(1)-SV 1.60 0.00 1.27 0.00 1.39 0.00 1.27 0.00 1.07 0.00 0.94 0.00 0.88 0.00 1.15
+BVAR(1)-CSV 0.73 0.18* 0.77 0.18* 0.82 0.17* 0.90 0.18* 0.93 0.18* 0.94 0.18* 0.91 0.18* 0.88
+DNS 1.15 0.00 1.10 0.00 1.05 0.00 1.08 0.00 1.11 0.00 1.13 0.00 1.04 0.00 1.09
+TVP-QNS 0.76 1.00** 0.75 1.00** 0.73 1.00** 0.76 1.00** 0.76 1.00** 0.74 1.00** 0.64 1.00** 0.73
+36 Month-Ahead Forecast
+TVP-AR(1)-SV 0.00 0.00 0.00 0.00 0.00 0.00 2.84 0.00
+BVAR(1)-CSV 1.13 0.00 1.17 0.00 1.19 0.00 1.24 0.00 1.26 0.00 1.27 0.00 1.25 0.00 1.23
+DNS 1.04 0.00 1.03 0.00 1.01 0.00 1.05 0.00 1.08 0.00 1.12 0.00 1.10 0.00 1.07
+TVP-QNS 0.79 1.00** 0.79 1.00** 0.79 1.00** 0.84 1.00** 0.86 1.00** 0.87 1.00** 0.79 1.00** 0.83
+Notes: Superscripts and ∗∗ indicate inclusion in the model confidence sets $\widehat{\mathcal{M}_{90\%}^{*}}$ and $\widehat{\mathcal{M}_{75\%}^{*}}$ respectively, as per Hansen et al.
+(2011).
+The reported average score is computed across all 17 maturities.
+A dash (’–’) indicates that the quantile scores for the TVP-AR(1)-SV model are exceptionally large and thus omitted from the table for clarity.
+Tables 2 and 3 present the 10th and 90th percentile quantile scores (QS) from the out-of-sample forecasting exercise for selected maturities.
+We also report the corresponding p-values from the Model Confidence Set (MCS) procedure proposed by Hansen et al.
+(2011).
+The results suggest that the proposed quantile Nelson-Siegel model delivers superior tail risk forecasts, particularly at longer forecast horizons and for long-term maturities.
+At shorter horizons, however, the BVAR(1)-CSV model remains competitive and, in some cases, outperforms the proposed model in forecasting tail risks.
+Importantly, the quantile Nelson-Siegel model consistently outperforms its conditional mean counterparts, based on Diebold et al.
+(2006), across all forecast horizons and quantile levels.
+Moreover, the MCS analysis shows that the proposed model is always included in the $\widehat{\mathcal{M}_{90\%}^{*}}$ and $\widehat{\mathcal{M}_{75\%}^{*}}$ confidence sets, especially for longer maturities and extended forecast horizons.
+By contrast, the three conditional mean models are frequently excluded from these sets.
+Overall, these findings provide strong empirical support for the proposed time-varying parameter quantile Nelson-Siegel model as a robust tool for forecasting tail risk.
+Table 3: Average 90% (Upper-Tail) Quantile Scores for Selected Maturities Relative to an AR(1) Benchmark
+Model 3m $p_{\text{MCS}}$ 6m $p_{\text{MCS}}$ 12m $p_{\text{MCS}}$ 24m $p_{\text{MCS}}$ 36m $p_{\text{MCS}}$ 60m $p_{\text{MCS}}$ 120m $p_{\text{MCS}}$ Avg.
+1 Month-Ahead Forecast
+TVP-AR(1)-SV 0.37 0.00 0.39 0.01 0.49 0.03 0.65 0.00 0.75 0.00 0.85 0.00 0.90 0.00 0.69
+BVAR(1)-CSV 0.21 1.00** 0.24 1.00** 0.31 1.00** 0.40 1.00** 0.46 1.00** 0.54 1.00** 0.62 1.00** 0.44
+DNS 1.09 0.00 1.04 0.00 1.15 0.00 1.16 0.00 1.12 0.00 1.09 0.00 1.55 0.00 1.17
+TVP-QNS 0.53 0.00 0.57 0.00 0.63 0.04 0.59 0.00 0.56 0.02 0.56 0.70** 0.83 0.00 0.61
+12 Month-Ahead Forecast
+TVP-AR(1)-SV 0.89 1.00** 0.85 1.00** 0.90 1.00** 0.92 1.00** 0.91 1.00** 0.90 1.00** 0.87 1.00** 0.89
+BVAR(1)-CSV 0.83 0.93** 0.87 0.93** 0.93 0.93** 0.97 0.93** 0.99 0.94** 1.03 0.93** 1.09 0.94** 0.98
+DNS 1.07 0.00 1.08 0.00 1.10 0.00 1.07 0.00 1.03 0.00 0.98 0.00 1.06 0.00 1.04
+TVP-QNS 0.99 0.93** 0.99 0.93** 0.98 0.93** 0.90 0.93** 0.83 0.94** 0.76 0.93** 0.77 0.94** 0.86
+36 Month-Ahead Forecast
+TVP-AR(1)-SV 0.00 0.00 0.00 0.00 0.00 0.00 1.37 0.00
+BVAR(1)-CSV 1.68 0.00 1.68 0.00 1.70 0.00 1.72 0.00 1.71 0.00 1.76 0.00 1.90 0.00 1.75
+DNS 1.07 0.05 1.06 0.26∗∗ 1.06 0.74∗∗ 1.03 0.05 0.99 0.00 0.96 0.00 1.03 0.00 1.02
+TVP-QNS 1.06 1.00** 1.03 1.00** 1.01 1.00** 0.95 1.00** 0.90 1.00** 0.86 1.00** 0.87 1.00** 0.93
+Notes: Superscripts and ∗∗ indicate inclusion in the model confidence sets $\widehat{\mathcal{M}_{90\%}^{*}}$ and $\widehat{\mathcal{M}_{75\%}^{*}}$ respectively, as per Hansen et al.
+(2011).
+The reported average score is computed across all 17 maturities.
+A dash (’–’) indicates that the quantile scores for the TVP-AR(1)-SV model are exceptionally large and thus omitted from the table for clarity.
 4.3 In-Sample Results
@@
 This section presents our empirical in-sample results.
-The first subsection details the posterior estimates of yield curve factors for each quantile level.
-The second subsection analyses the distributional features of short- and long-term interest rates during recent crises (the Great Recession and the COVID-19 pandemic).
-Lastly, the third subsection investigates the interaction between yield curve factors and the macroeconomy from a quantile perspective.
+The first subsection reports posterior estimates of the yield curve factors and the decay parameter across various quantile levels.
+The second subsection investigates the distributional properties of short- and long-term interest rates, as well as selected Treasury maturities, during recent crises—the Great Recession and the COVID-19 pandemic.
+Finally, the third subsection analyses the relationship between yield curve factors and macroeconomic variables from a quantile perspective.
+4.3.1 Posterior Estimates of the Decay Parameter
+Figure 2: Posterior distribution of $\lambda(\tau)$ for the 10%, 25%, 50%, 75% and 90% quantiles, with the posterior mean (dashed line).
+Figure 2 presents the posterior distribution of the decay parameter $\lambda(\tau)$ across five selected quantiles (10%, 25%, 50%, 75%, and 90%).
+The results indicate that, for each quantile, the posterior mass of $\lambda(\tau)$ is concentrated in the range of approximately 0.06 to 0.07.
+This is broadly consistent with the fixed value of $\lambda=0.0609$ proposed by Diebold et al.
+(2006).
 4.3.2 Quantile Yield Curve Factor Estimates
-Figures 2, 3, and 4 depict the posterior mean estimates of time-varying yield curve factors at five quantiles (10%, 25%, 50%, 75%, 90%), providing insights into their dynamic evolution.
+Figures 3, 4, and 5 depict the posterior mean estimates of time-varying yield curve factors at five quantile levels, providing insights into their dynamic evolution.
 A consistent finding across all graphs is the presence of substantial temporal variation.
@@
 Figure 3: Posterior mean estimates of the level factor for each of the five quantiles (10%, 25%, 50%, 75%, 90%). NBER recession periods in grey shades.
-Turning to the slope and curvature of the yield curve in Figures 3 and 4, the degree of uncertainty associated with these factors varies over time.
+Turning to the slope and curvature of the yield curve in Figures 4 and 5, the degree of uncertainty associated with these factors varies over time.
 Notably, the curvature factor exhibits higher volatility across all quantiles than its slope counterpart.
-[^1] Notice that crossing of the estimated conditional quantiles for the curvature factor is not harmful since it does not have a direct link to one particular yield.
+[^1] We impose the monotonicity constraint specified in equation (18) on the short-term interest rate, defined as the sum of the level and slope factors.
+Figure 10 in the appendix displays the posterior estimates of the short-term rate across the five quantiles, which exhibit no crossing and are thus consistent with the imposed constraint.
+However, this restriction does not imply monotonicity in the slope factor alone, which may still exhibit crossings across quantiles.
+In contrast, crossings in the conditional quantiles of the curvature factor are not of concern, as the curvature does not correspond directly to any specific yield maturity.
 A distinct pattern is evident before 1985, where the posterior estimates for the slope factor across all five quantiles are closely clustered, indicating a period of low uncertainty.
@@
 Conversely, the proposed quantile approach enables deeper and more detailed analyses of the term structure while maintaining a low computational cost.
-4.2.2 The Distributional Features of the Short- and Long-Term Interest Rate
-The proposed novel framework provides a distinct advantage by facilitating the investigation of distributional characteristics in both short- and long-term interest rates over time.
-As mentioned earlier, the long-term interest rate is identified as the level factor, while the short-term interest rate is a composite of the level and slope factors.
-We adhere to the methodology outlined by Mitchell et al.
-(2022) to construct the distributions of these interest rates.
-Employing an equal weighting scheme, we aggregate all MCMC draws across five quantiles and apply the Epanechnikov kernel, as proposed by Gaglianone and Lima (2012).
+4.3.3 Distributional Characteristics of Treasury Yields Across the Term Structure
+The proposed novel framework offers a distinct advantage by enabling the examination of the distributional properties of both short- and long-term interest rates, as well as selected yield maturities, over time.
+As previously discussed, the long-term interest rate is identified with the level factor, whereas the short-term interest rate reflects a combination of the level and slope factors.
+We follow the methodology outlined by Mitchell et al.
+(2024) to construct the empirical distributions of these rates and maturities.
+Specifically, we adopt an equal-weighting scheme to aggregate all MCMC draws across five quantiles and apply the Epanechnikov kernel, as proposed by Gaglianone and Lima (2012).
 The previous analysis revealed an increase in the degree of dispersion among quantiles during post-recessionary periods, particularly in the 2000s.
@@
 Figure 6: Distribution of the short-term and long-term (level) factors during the Great Recession (left) and the COVID-19 pandemic (right).
-Focusing on the short-term interest rate first, discernible evidence of a pronounced distributional shift is observed across the three selected months for both recessions.
-Notably, during the Great Recession (September 2009), a manifestation of multi-modality in the distribution is apparent.
-Additionally, after the COVID-19 recession (December 2020), the distribution became less symmetric and more negatively skewed to the left.
-Turning to the long-term interest rate, the evidence of distribution shifts is more prominent during the COVID-19 recession than during the Great Recession.
-The long-term interest rate distribution exhibits increased leftward skewness during the Great Recession (September 2009) and the subsequent period (December 2009).
-Conversely, during the COVID-19 recession, the distribution becomes more positively skewed to the right.
-Therefore, our results highlight that both short- and long-term interest rate distributions display heightened skewness and asymmetric shapes during recessionary periods.
-Thus highlighting the importance of using quantile models instead of standard Nelson-Siegel models that are not able to detect the left and right tails of the distribution.
-4.2.3 The Interaction between the Yield Curve and Macroeconomy - Quantile VAR Parameter estimates
-The proposed framework, as delineated in eq.
-(9), introduces an avenue for interaction between yield curve factors and macroeconomic variables by incorporating a VAR-like state equation.
-The estimated off-diagonal VAR coefficients for various quantile levels are presented in Figure 6, shedding light on the directional effects between yield curve factors and macroeconomic variables.
-Panel A
-Panel B
-Figure 6: Posterior mean and credible intervals for the VAR coefficients across different quantile levels.
-Panel A of the figure conveys the directional impact of yield curve factors on macroeconomic variables, while Panel B reports the converse scenario.
-The patterns of influence of each yield factor on the three macroeconomic variables exhibit remarkable similarity.
-Precisely, evidence of a U-shaped effect across quantiles is discerned for industrial production and inflation, with an inverse relationship observed for the Fed funds rate.
-Conversely, the impact of macroeconomic variables on the yield curve factor demonstrates higher heterogeneity.
-Industrial production exerts a significant positive effect on both level and curvature, whereas its impact on the slope is mixed.
-The effects of inflation exhibit qualitative similarity, albeit with a smaller magnitude on the level and a higher impact on the slope.
-In contrast, the Fed funds rate positively influences all yield curve factors.
-Analysing the relationship with quantile levels, the effects on the slope exhibit a U-shaped pattern for industrial production and inflation, while the Fed funds rate demonstrates an inverse U-shaped relationship.
-Meanwhile, the relationship with the level remains constant across quantiles.
-In summary, our findings provide substantial evidence that the interaction between yield curve factors and the macroeconomy significantly varies across quantiles.
-4.3 Forecasting Yields Tail Risk
-We conduct a pseudo-out-of-sample forecasting exercise to assess the utility of our novel quantile Nelson-Siegel approach, comparing it against the corresponding standard Nelson-Siegel conditional mean counterpart for forecasting tail risks in zero-coupon yields.
-The initial holdout period spans from November 1971 to January 2000, with a subsequent forecast evaluation period extending from February 2000 to December 2022 (comprising 264 evaluation periods).
-We generate 1st, 6th, and 12th step-ahead tail risk forecasts and employ the quantile score (QS) (see Giacomini and Komunjer, 2005; Carriero et al., 2022) as the forecast evaluation metric to gauge the performance of the two models.
-Following Gneiting and Ranjan (2011), we define the QS as:
-$QS_{\tau,i,t}=(y_{t}(m)-\mathcal{Q}_{\tau,i,t})-(\tau-\mathbb{I}\{y_{t}(m)\leq\mathcal{Q}_{\tau,i,t}\}).$ (17)
-Here, $QS_{\tau,i,t}$ denotes the predictive quantile of the $m$-th yield, and $\mathbb{I}\{y_{t}(m)\leq\mathcal{Q}_{\tau,i,t}\}$ takes the value 1 if the realised value is at or below the predictive quantile and 0 otherwise.
-We evaluate the QS in the upper and lower tails by setting $\tau=0.9$ and $\tau=0.1$, respectively.
-Table 1 presents the 10th and 90th percentiles of QS for the out-of-sample forecasting exercise.
-The top panel of Table 1 displays the average QS of our proposed approach relative to their corresponding conditional mean counterparts for each of the 17 yields across the evaluation period.
-A quantile score less than 1 implies that our proposed approach outperforms the corresponding conditional mean counterpart.
-Notably, our proposed approach consistently produces superior tail forecasts for all 17 yields and across all forecast horizons compared to their corresponding conditional mean counterparts.
-We further assess the significance of our forecasting results through a Diebold-Mariano test, with results presented in the bottom panel of Table 1, clearly indicating the statistical significance of our forecasting results in favour of our proposed approach.
-Additionally, we depict the average raw QS across the 17 yields for both models over time in Figure 7.
-The top and bottom panels show the 10th and 90th percentiles of QS, respectively.
-Three key features emerge from this figure.
-First, reinforcing the results in Table 1, our proposed quantile Nelson-Siegel model consistently outperforms their corresponding conditional mean counterparts across all horizons and quantile levels, with the evidence being stronger for 1-step ahead forecasts.
-Second, the 1-step ahead QS (for both 10th and 90th percentiles) for the conditional mean model exhibits significantly more instability over time than our quantile approach.
-Finally, we find strong evidence in favour of the proposed time-varying parameters quantile Nelson-Siegel model for forecasting the left tail of the yield distribution, suggesting its ability to more appropriately capture the lower bound of the yield curve compared to standard conventional models.
-Figure 7: Quantile Score over time for the 1-, 6-, and 12-steps ahead forecast horizons, at quantile levels 10% (top plots) and 90% (bottom plots), for the QR-NS model (blue) and the standard NS model (red).
-Shaded areas represent NBER recession periods.
-Table 1: Forecasting results for the quantile NS relative to standard NS model, together with Diebold-Mariano test. Out-of-sample evaluation period: 2000 - 2022.
-Ratio of quantile scores (quantile NS over standard NS)
-10% QS
-Horizon 3 m 6 m 9 m 12 m 15 m 18 m 21 m 24 m 30 m 36 m 48 m 60 m 72 m 84 m 96 m 108 m 120 m Average
-1-step ahead 0.28 0.34 0.39 0.43 0.44 0.45 0.45 0.45 0.43 0.42 0.41 0.42 0.45 0.50 0.54 0.60 0.65 0.45
-6-step ahead 0.61 0.63 0.64 0.65 0.65 0.65 0.64 0.64 0.63 0.61 0.59 0.58 0.57 0.57 0.57 0.58 0.59 0.61
-12-step ahead 0.67 0.68 0.69 0.70 0.70 0.71 0.71 0.70 0.70 0.69 0.67 0.66 0.65 0.63 0.63 0.62 0.62 0.67
-90% QS
-Horizon 3 m 6 m 9 m 12 m 15 m 18 m 21 m 24 m 30 m 36 m 48 m 60 m 72 m 84 m 96 m 108 m 120 m Average
-1-step ahead 0.47 0.55 0.56 0.55 0.53 0.51 0.50 0.49 0.49 0.49 0.50 0.53 0.54 0.56 0.57 0.57 0.56 0.53
-6-step ahead 0.98 0.98 0.96 0.93 0.90 0.88 0.86 0.84 0.80 0.77 0.73 0.71 0.69 0.69 0.69 0.69 0.68 0.81
-12-step ahead 0.96 0.96 0.95 0.93 0.92 0.91 0.89 0.88 0.86 0.84 0.82 0.80 0.79 0.78 0.77 0.76 0.75 0.86
-Diebold-Mariano test, significance level: 1 (10%), 2 (5%), 3 (1%)
-10% QS
-Horizon 3 m 6 m 9 m 12 m 15 m 18 m 21 m 24 m 30 m 36 m 48 m 60 m 72 m 84 m 96 m 108 m 120 m Average
-1-step ahead 3.00 3.00 3.00 3.00 3.00 3.00 3.00 3.00 3.00 3.00 3.00 3.00 3.00 3.00 3.00 3.00 3.00 3.00
-6-step ahead 3.00 3.00 3.00 3.00 3.00 3.00 3.00 3.00 3.00 3.00 3.00 3.00 3.00 3.00 3.00 3.00 3.00 3.00
-12-step ahead 3.00 3.00 3.00 3.00 3.00 3.00 3.00 3.00 3.00 3.00 3.00 3.00 3.00 3.00 3.00 3.00 3.00 3.00
-90% QS
-Horizon 3 m 6 m 9 m 12 m 15 m 18 m 21 m 24 m 30 m 36 m 48 m 60 m 72 m 84 m 96 m 108 m 120 m Average
-1-step ahead 3.00 3.00 3.00 3.00 3.00 3.00 3.00 3.00 3.00 3.00 3.00 3.00 3.00 3.00 3.00 3.00 3.00 3.00
-6-step ahead 0.00 0.00 2.00 3.00 3.00 3.00 3.00 3.00 3.00 3.00 3.00 3.00 3.00 3.00 3.00 3.00 3.00 2.59
-12-step ahead 1.00 1.00 2.00 3.00 3.00 3.00 3.00 3.00 3.00 3.00 3.00 3.00 3.00 3.00 3.00 3.00 3.00 2.71
+Focusing first on the short-term interest rate, we observe clear evidence of a pronounced distributional shift across the three selected months during both the Great Recession and the COVID-19 recession.
+In particular, following the COVID-19 recession (December 2020), the distribution of the short-term rate becomes less symmetric and exhibits increased negative skewness.
+Turning to the long-term interest rate, distributional changes appear more pronounced during the COVID-19 recession than during the Great Recession.
+After the Great Recession (December 2009), the distribution becomes more negatively skewed.
+In contrast, during the COVID-19 recession, the distribution shifts towards positive skewness, indicating a different form of market adjustment at the long end of the yield curve.
+Figure 7 presents the estimated conditional distributions of three selected Treasury maturities—3 months, 3 years, and 10 years—representing the short, intermediate, and long segments of the yield curve, respectively, across the two crisis periods.
+Prior to the Great Recession, the distributions of these maturities are relatively similar.
+However, in the aftermath of the Great Recession, we observe a clear location shift across all maturities, with the 10-year maturity displaying increased left-skewness.
+During the COVID-19 recession, only the long end of the yield curve undergoes a location shift, accompanied by greater dispersion.
+Following the COVID-19 episode, all three maturities exhibit a location shift, with the short and intermediate maturities also displaying increased negative skewness.
+These results indicate that short- and long-term interest rates, as well as selected yield maturities, experience heightened skewness and asymmetry during recessionary periods.
+This highlights the empirical relevance of quantile-based approaches, which allow for a more flexible representation of the conditional distribution—particularly in the tails—than traditional factor models such as Nelson-Siegel, which impose symmetry and are limited in their ability to capture such nonlinearities.
+Figure 7: Distribution of the 3-month, 3-year, and 10-year Treasury yields during the Great Recession (top panel) and the COVID-19 pandemic period (bottom panel).
+The solid blue line corresponds to the 3-month maturity, the red dashed line represents the 3-year maturity, and the black dashed line denotes the 10-year maturity.
+4.3.4 The Interaction between the Yield Curve and Macroeconomy
+The proposed framework, as specified in Equation (9), enables an in-depth analysis of the dynamic interaction between yield curve factors and macroeconomic variables by incorporating a VAR-like state equation.
+To explore these interactions, we compute the generalized forecast error variance decomposition (GFEVD) across five quantiles, evaluating the contribution of shocks to yield curve factors on macroeconomic variables—and vice versa—as illustrated in Figures 8 and 9, respectively.
+We begin by examining the effects of yield curve shocks on macroeconomic variables (Figure 8).
+The analysis reveals that curvature shocks generate heterogeneous impacts across the distribution of macroeconomic outcomes.
+Notably, these shocks have a more pronounced effect on the upper tail of all three macroeconomic variables—particularly at longer forecast horizons—indicating heightened sensitivity in more expansionary states of the economy.
+In contrast, level shocks increasingly influence the lower tail of industrial production over time, suggesting growing downside risks in response to shifts in the overall level of the yield curve.
+We next assess the influence of macroeconomic shocks on the yield curve (Figure 9).
+Shocks to industrial production significantly affect the lower tails of all three yield curve factors, implying that negative real activity shocks disproportionately compress the yield distribution.
+By contrast, inflation shocks primarily impact the upper tails of the slope and curvature factors, with these effects intensifying at longer horizons—consistent with rising term premia during inflationary episodes.
+Additionally, federal funds rate shocks increasingly affect the upper tail of the level factor over time, underscoring the asymmetric transmission of monetary policy to the term structure.
+Figure 8: Posterior estimates of the generalized forecast error variance decomposition for the macroeconomic variables (columns) on Yield Curve Factors (rows).
+The colored bars represent different quantiles: blue (10%), orange (25%), yellow (50%), purple (75%), and green (90%).
+Figure 9: Posterior estimates of the generalized forecast error variance decomposition for the Yield Curve Factors (columns) on Yield Curve Factors (rows).
+The colored bars represent different quantiles: blue (10%), orange (25%), yellow (50%), purple (75%), and green (90%).
+In summary, these results provide robust evidence that the interaction between yield curve factors and macroeconomic variables is highly dependent on the quantile considered.
+This highlights the importance of employing quantile-based approaches, which can capture distributional asymmetries and tail behavior that standard linear models may overlook.
 5 Conclusions
-In this study, we introduce a novel time-varying parameters quantile Nelson-Siegel (TVP-QR-NS) model within a Bayesian framework and apply it to an empirical analysis of the US term structure of interest rates.
-Our in-sample findings reveal an increased dispersion in yield curve factors across quantiles in the post-recession period.
-Additionally, temporal variations in the distribution of short- and long-term interest rates are observed during the two most recent recessions.
-Furthermore, evidence indicates significant fluctuations in the interaction between yield curve factors and the macroeconomy across quantiles.
-Finally, through an out-of-sample forecasting exercise, we demonstrate the superior performance of our proposed approach in tail risk forecasting across all yields compared to its standard conditional mean counterpart.
+This study introduces a novel time-varying parameter quantile Nelson-Siegel (TVP-QR-NS) model, estimated within a Bayesian framework, to analyse the US term structure of interest rates.
+The out-of-sample forecasting results demonstrate that the TVP-QR-NS model consistently outperforms conventional benchmarks in forecasting tail risks, particularly at longer horizons and for long-term maturities.
+Importantly, it is the only model consistently included in the model confidence sets for these maturities and horizons, while standard conditional mean models are frequently excluded.
+The in-sample analysis reveals significant dispersion in yield curve factors across quantiles, especially following recessionary periods.
+The analysis also uncovers pronounced shifts in the distributions of short- and long-term interest rates—as well as yields at selected maturities—during the Great Recession and the COVID-19 pandemic, highlighting the importance of capturing variation across both time and quantiles.
+In addition, strong evidence is found of quantile-dependent dynamics in the relationship between yield curve factors and macroeconomic conditions.
+Overall, these findings highlight the robustness and practical relevance of the TVP-QR-NS framework for risk management and macro-financial policy analysis.
+Appendix A Appendix
+A.1 Sampling the Decay Parameter
+As mentioned above, it has so far been common practice to fix $\lambda$ to be a particular value.
+On the other hand, we are agnostic and estimate $\lambda(\tau)$ given the data for each quantile.
+Given a uniform prior for $\lambda(\tau)$, $p(\lambda(\tau))\sim U(a,b)$, the conditional posterior of $(\lambda(\tau)|\bullet)$ is given by
+$(\lambda(\tau)|\bullet)\propto p(\mathbf{Y},\mathbf{w}|\bm{\beta})p(\lambda(\tau)),$
+where $a<\lambda(\tau)<b$.
+Since the support of this conditional density is bounded and non-standard, we draw from this conditional density using a Griddy-Gibbs step (Ritter and Tanner, 1992).
+This is done using the following steps:
+  1. 1.
+    Construct a grid with grid points $\hat{\lambda(\tau)}_{1},\ldots,\hat{\lambda(\tau)}_{R}$, where $\hat{\lambda(\tau)}_{1}=a$ and $\hat{\lambda(\tau)}_{R}=b$.
+  2. 2.
+    Compute $F_{i}=\sum_{j=1}^{i}p(\hat{\lambda(\tau)}_{j}|\bullet)$.
+  3. 3.
+    Generate $U$ from a standard uniform distribution.
+  4. 4.
+    Find the smallest positive integer $k$ such that $F_{k}\geqslant U$ and return $\lambda(\tau)=\hat{\lambda(\tau)}_{k}$.
+A.2 Additional In-Sample Results
+Figure 10: Posterior mean estimates of the short-term rate (Level + Slope) for each of the five quantiles (10%, 25%, 50%, 75%, 90%). NBER recession periods in grey shades.